# On the Low-Temperature MCMC threshold:

### v2

@@ Abstract.
Over the last years, there has been a significant amount of work studying the power of specific classes of computationally efficient estimators for multiple statistical parametric estimation tasks, including the estimator classes of low-degree polynomials, spectral methods, and others. Despite that, our understanding of the important class of MCMC methods remains quite poorly understood. For instance, for many models of interest, the performance of even zero-temperature (greedy-like) MCMC methods that simply maximize the posterior remains elusive.
In this work, we provide an easy to check condition under which the low-temperature Metropolis chain maximizes the posterior in polynomial-time with high probability. The result is generally applicable, and in this work, we use it to derive positive MCMC results for two classical sparse estimation tasks: the sparse tensor PCA model and sparse regression. Interestingly, in both cases, we also leverage the Overlap Gap Property framework for inference (Gamarnik, Zadik AoS ’22) to prove that our results are tight: no low-temperature local MCMC method can achieve better performance. In particular, our work identifies the “low-temperature (local) MCMC threshold” for both sparse models.
Notably, in the sparse tensor PCA model our results indicate that low-temperature local MCMC methods significantly underperform compared to other studied time-efficient methods, such as the class of low-degree polynomials. Moreover, the “amount” by which they underperform appears connected to the size of the computational-statistical gap of the model. Specifically, an intriguing mathematical pattern is observed in sparse tensor PCA, which we call “the geometric rule”: the information-theory threshold, the threshold for low-degree polynomials and the MCMC threshold form a geometric progression for all values of the sparsity level and tensor power.
@@ 1. Introduction
Over the recent years, it has been revealed that a plethora of statistical estimation models exhibit computational-statistical gaps, i.e., parameter regimes for which some estimator can achieve small error, but no such polynomial-time estimator is known to achieve similar error guarantees. It is an intriguing and rapidly growing research direction to understand whether these gaps are “inherent” or an inability of our current polynomial-time statistical estimation techniques. Given that the (simpler [^1] Often the statistical models of interest include random noise assumptions, hence cannot be treated as “classic” computational questions about tasks with a worst-case input, such as e.g., the traveling salesman problem over a worst-case graph. It is this difference that makes statements about the classes such as $\mathcal{P}$ and $\mathcal{NP}$ not directly applicable to these statistical estimation settings.) worst-case question of $\mathcal{P}\not=\mathcal{NP}$ remains open in computational complexity theory, unfortunately proving that all polynomial-time estimators fail to close any given computational-statistical gap appears well beyond our current mathematical abilities.
@@
Despite that, over the recent years, researchers have constructed several tools to provide “rigorous evidence” for a computational-statistical gap of interest. Some of the profound approaches include average-case reductions to “hard regimes” of other statistical problems (see e.g., [BR13, MW15, BB20]), statistical physics inspired “hardness” heuristics [GMZ22], and also proofs that natural restricted classes of polynomial-time estimators fail to close these gaps. Some of the most popular studied such restricted classes are MCMC methods [Jer92, CMZ23], low-degree polynomials [SW22], semidefinite programs [BHK+19], message passing procedures [FVRS22], and the class of statistical query methods [FGR+17].
@@
Among these restricted classes of estimators perhaps the most recent attention has been received by the class of estimators that take the form of a low-degree (e.g., $O(1)$ or $O(\log n)$ degree) polynomial function of the data (first used in the context of hypothesis testing in [Hop18] and for estimation settings in [SW22]). A series of papers have built interesting tools to understand the exact threshold (in terms of some critical natural quantity of the model, often the sample size or signal to noise ratio) for a given statistical problem above which some low-degree polynomial works, and below which all low-degree polynomials fail (see e.g., [KWB19] for a survey). On top of that, a conjecture has been put forward in the context of hypothesis testing [Hop18] that – under mild noisy and symmetry assumptions [HW20, ZSWB22] – the threshold below which all $O(\log n)$-degree polynomials fail should in fact coincide with the threshold for which all polynomial-time methods fail as well. In other words, subject to this “low-degree conjecture”, in many classical statistical models the performance of $O(\log n)$-degree polynomials (which one can often directly analyze) allow us to characterize the performance of all polynomial-time estimators. Given the recent rapid growth of the low-degree literature for estimation problems following [SW22] (e.g., [MW22, MWZ23, LG23]), it is expected that a similar low-degree conjecture will be put forward in the near future in terms of statistical estimation problems as well.
@@
Despite this recent popularity of analyzing the performance of low-degree polynomials for statistical estimation settings, it is natural to wonder what happens with other natural classes of polynomial-time estimators that are significantly different from low-degree polynomials (e.g., estimators that are not even well-approximated by them). One such highly canonical class of statistical estimators (both in theory and applications of statistics) is the class of Markov Chain Monte Carlo (MCMC) methods that try to sample from the posterior (or slight variants of it), such as the Metropolis chain in discrete time and Langevin dynamics in continuous time. Researchers have studied this important class for a large family of models, but often arrive at significantly less satisfying theoretical results than the ones for the literature of low-degree methods. Among these MCMC results, most of them concern lower bounds, that is provable failures of classes of MCMC method to achieve a certain guarantee. Some influential such results include MCMC lower bounds for the planted clique problem [Jer92, GZ19, CMZ23], sparse principal component analysis (PCA) [GJS21, AWZ23] and tensor PCA [BAGJ20]. Importantly, many of the recent such results leverage a suggested proof framework known as Overlap Gap Property for inference to obtain bottlenecks for the studied Markov chains, first introduced in [GZ22a]. On the other hand, there are significantly fewer positive MCMC results known for such estimation problems, including though interesting positive results about Bayesian variable selection [YWJ16], tensor PCA [BAGJ20], and recently also the planted clique model [GJX23] and the stochastic block model [LMR+24]. Yet, even concerning these relatively few positive results, most of them are an outcome of a technical detailed analysis tailored to the specific model of interest. In particular, as opposed to lower bounds and the proof framework of [GZ22a], there is a lack of a general “easy-to-use” strategy of how can one obtain a positive MCMC result for a new model of interest, besides directly trying to bound the mixing time (or analyze it’s locally stationary measures as in the recent work [LMR+24]). A significant motivation for this work is to answer the following question:
@@
Besides only obtaining positive MCMC results, it is of course highly desirable to obtain the “optimal” MCMC results, in the sense of locating exactly what is the minimal sample size (or signal to noise ratio) so that an element of some natural class of MCMC methods succeeds. Unfortunately, it is significantly understudied how one could locate exactly the “MCMC threshold” for a generic class of estimation settings, in sharp contrast with the generally applicable tools designed for the class of low-degree polynomials [SW22, YZZ24]. On top of that, the pursuit of the MCMC threshold in estimation settings is even more motivated because of recent surprising discoveries in the context of tensor PCA [BAGJ20] and planted clique [CMZ23]. The reason is that in both of these models the cited works have proven that a severe underperformance of certain natural MCMC methods takes place compared to other polynomial-time methods (something also noticed in the statistical physics literature [MKUZ19]). To make it more precise, let us focus on the tensor PCA case where for some integer $t\geq 2$ one observes the $t$-tensor $Y=\lambda v^{\otimes t}+W$ for some $v\in S^{d-1}$ chosen uniformly at random from the sphere and $W$ has i.i.d. $N(0,1)$ entries. The task is to recover $v$ by observing $Y$. While the minimum $\lambda=\lambda_{d}>0$ for which this is possible by some (time-unconstrained) method scales like $\lambda_{\mathrm{STATS}}\approx\sqrt{d}$, polynomial-time methods are known to begin to work only when $\lambda_{\mathrm{ALG}}\approx d^{t/4}$ and below $\lambda_{\mathrm{ALG}}$ all $O(\log n)$-degree polynomials are known to fail, suggesting that $\lambda_{\mathrm{ALG}}$ is the computational threshold for this setting. Yet, if one runs Langevin dynamics that tries to sample from the (slightly temperature misparametrized) posterior, [BAGJ20] proves that the dynamics can recover $v$ in polynomial-time at the threshold $\lambda_{\mathrm{MCMC}}\approx d^{t/2-1/2}.$ Notice that for $t\geq 3$, $d^{t/4}=o(d^{t/2-1/2})$. It is perhaps striking the underperformance by a polynomial-in-$d$ factor of the Langevin dynamics for this seemingly simple task. For such reasons, this phenomenon that tensor PCA undergoes is known in the community as a “local-to-computational gap”. A partial motivation of this work is to understand how generic the existence of local-to-computational gaps is in statistical estimation settings.
@@
Besides the mathematical elegance, a positive answer to such a question would mean that for a class of models, as long as there is a computational statistical trade-off, i.e., $\lambda_{\mathrm{ALG}}/\lambda_{\mathrm{STATS}}=\omega(1)$ then there is a local-to-computational gap as well, i.e., $\lambda_{\mathrm{MCMC}}/\lambda_{\mathrm{ALG}}=\omega(1)$. In this work, we interestingly reveal that the geometric rule does indeed hold beyond just the tensor PCA model, and in particular for the sparse matrix and tensor PCA models as well.
@@ 1.1. Contributions
We study the parameter estimation set-up, where for some parameter $\theta^{*}\in\Theta\subseteq\mathbb{R}^{d}$ sampled from some prior $\mathcal{P}$ we observe a dataset $\mathcal{D}\sim\mathbb{P}_{\theta^{*}}$. We assume, for simplicity, that $\Theta$ is a finite set of parameters and the prior is fully supported on $\Theta$. The goal of the statistician is to exactly recover the value of $\theta^{*}$ from observing $\mathcal{D}$, with high probability over both the randomness of the prior and the channel as $d$ grows to infinity. Here and everywhere in the paper, by saying that an event holds with high probability as $d$ grows to infinity, we mean that it’s probability converges to one as $d$ grows to infinity.
@@
For convenience in what follows, we refer to such a family of local Markov chains as “low-temperature MCMC methods”. We are interested under which assumptions low temperature MCMC methods can recover $\theta^{*}$ with high probability.
@@ 1.1.1. A general positive result
As we mentioned in the Introduction, the positive results for such low temperature MCMC methods are few, often quite technical, and usually an outcome of an analysis crucially tailored to each case of interest.
Our first contribution is a general positive “black-box” result (Theorem 2.3) which offers a simple condition under which some low-temperature MCMC method succeeds in polynomial-time to optimize the posterior, and in particular recover $\theta^{*}$ with high probability. Our result in fact applies for any optimization problem over a discrete feasible set and we expect to be of independent interest to the discrete optimization community as well.
@@
As long as solely this condition is satisfied, we prove that simply running the (low-temperature) Metropolis chain on this neighborhood graph for polynomial time is guaranteed to find the optimizer of the posterior $\theta^{*}$ with high probability. As we mentioned above, to the best of our knowledge, no similar result is known even in the discrete/combinatorial optimization community. The proof is simple and is based on a general canonical paths construction [Jer03] and could be potentially of independent interest.
To show the generality of Theorem 2.3 we derive two positive MCMC results for two of the most classical, but significantly different, sparse estimation problems with well-studied computational-statistical gaps. First, we focus on the sparse tensor PCA model (also called, constant high order clustering) [LZ22, Cd21] which is a special case of a Gaussian additive model, i.e., one observes a sparse signal drawn from some prior where Gaussian noise is added to it. Second, we also apply it to Gaussian sparse regression [GZ22a] which is a special case of a Linear Model, i.e., one observes i.i.d. samples of a (random) linear transformation of a sparse signal where Gaussian noise is added to each sample. Interestingly, using the Overlap Gap Property framework from [GZ22a] we can also show that both the obtained low-temperature MCMC results are tight, in the sense that a large class of low-temperature MCMC results fail to work in polynomial-time, whenever the conditions of our positive result do not hold. In particular, these results allow us to characterize the exact power of a large class of low-temperature MCMC methods for both of these settings.
@@ 1.1.2. Sparse Tensor PCA
The sparse tensor PCA model we focus on is a canonical Gaussian principal component analysis setting with a large literature in theoretical statistics (see e.g., [BI13, BIS15, CLR17, XZ19, NWZ20, LZ22] and references therein). Moreover, the setting is greatly motivated by a wide range of applications in many domains, such as genetics, social sciences, and engineering (see e.g., [HM18] for a survey). More specific interesting examples include topics such as microbiome studies [FSI+12] and multi-tissue multi-individual gene expression data [WFS19].
Now, from the theoretical standpoint it is known [LZ22, NWZ20] that the optimal estimator (MAP) can recover $v$ if and only if $\lambda\geq\lambda_{\mathrm{STATS}}(t,k)$ for some
@@
Moreover, based on average-case reductions to hypergraphs variants of the planted clique problem [LZ22], polynomial-time estimators are expected to work if and only if $\lambda\geq\lambda_{\mathrm{ALG}}(t,k)$ for some
@@
It should be noted that the mentioned algorithmic lower bound when $\lambda=o(\lambda_{\mathrm{ALG}}(t,k))$ from [LZ22] applies in principle for a slight variant of the model of sparse tensor PCA compared to one we introduced above, which they denote by $\mathrm{CHC}_{R}$. In $\mathrm{CHC}_{R}$ one observes $Y=\lambda v_{1}\otimes v_{2}\otimes\ldots\otimes v_{t}+W$ for $v_{i},i=1,\ldots,t$ which are independent $k$-sparse binary vectors chosen uniformly at random. Note that in our case we focus instead on the case $v_{1}=\ldots=v_{t}=v$ where $v$ is a $k$-sparse binary vectors chosen uniformly at random. Often in the literature, the model we study is called the symmetric version of a tensor PCA model, while the one studied in [LZ22] is called the asymmetric version (see e.g., the discussion in [DH21]).
Yet, a quick check in the proof of their lower bound [LZ22, Theorem 16] shows that in fact the authors first prove the conditional failure of polynomial-time algorithms when $\lambda=o(\lambda_{\mathrm{ALG}}(t,k))$ for our case of interest, that is when $v_{1}=\ldots=v_{t}$, and then add an extra step to also prove it for the case that $v_{i},i=1,\ldots,t$ are independent. In particular, their result applies almost identically to the case of symmetric sparse tensor PCA we focus on this work. For reasons of completeness, in Section 11, we describe exactly the slightly modified argument from [LZ22] and how it applies in our symmetric setting.
A careful application of our “black box” positive MCMC result interestingly implies that for all $t\geq 2$ and various values of the sparsity level $\omega(1)=k=o(p),$ if $\beta>0$ is large enough, the Metropolis chain with stationary measure $\pi_{\beta}$ mixes in polynomial time as long as $\lambda\geq\lambda_{\mathrm{MCMC}}(t,k)$ for a significantly different threshold compared to $\lambda_{\mathrm{ALG}}(t,k)$ given by
@@
See Theorem 3.3 for the exact statement.
Satisfyingly, we also prove that when $\lambda=\tilde{o}(\lambda_{\mathrm{MCMC}}(t,k))$ a negative result for all low temperature reversible local MCMC methods (including the Metropolis chain) holds. Specifically, we prove in Theorem 3.6 that for large enough $\beta>0$ all local reversible Markov chains with stationary measure $\pi_{\beta}$ mix in super-polynomial time. Notably, our result significantly improves upon the prior work [AWZ23] on sparse PCA where a similar negative result for local Markov chains was proven but only when $t=2$ and importantly under the more restrictive condition $\lambda=o(\lambda_{\mathrm{ALG}}(2,k)).$ The proof of Theorem 3.6 is the most technical contribution of the present work. We prove – via a delicate conditional second moment method – a bottleneck for the MCMC dynamics based on the Overlap Gap Property framework from [GZ22a]. The proof carefully leverages the “flatness” idea from random graph theory [BBSV19, GZ19], but adjusted to high dimensional Gaussian tensors, and could be of independent interest.
@@
We remark that this is a significant departure compared to the established tensor PCA result [BAGJ20] as described in (1.1), as it holds firstly under the sparsity assumption, and also it importantly applies even to the matrix case $t=2$ where tensor PCA does not exhibit either a computational-statistical gap or local-to-computational gap. Albeit interesting in its own right, the fact that the result applies when $t=2$ allows us also to see the geometric rule in simulations for relatively small values of $p,k$, which we find particularly satisfying (see Section 5). We suspect that a generalization of the geometric rule to a large family of Gaussian additive models is possible. The exact reason such an elegant formula appears in that generality remains an intriguing question for future work.

### v3

@@ Abstract.
Over the last years, there has been a significant amount of work studying the power of specific classes of computationally efficient estimators for multiple statistical parametric estimation tasks, including the estimator classes of low-degree polynomials, spectral methods, and others. Despite that, our understanding of the important class of MCMC methods remains quite poorly understood. For instance, for many simple statistical models of interest, the performance of even zero-temperature (randomized greedy-like) MCMC methods that simply maximize the posterior remains elusive.
In this work, we provide a structural geometric condition under which the low-temperature Metropolis chain maximizes the posterior in polynomial-time with high probability. The result is generally applicable, and in this work, we use it to derive positive MCMC results for two classical sparse estimation tasks: the sparse tensor PCA model and sparse regression. Interestingly, in both cases, we also carefully apply the Overlap Gap Property framework for inference (Gamarnik, Zadik AoS ’22) to prove that our results are tight: no low-temperature local MCMC method can achieve better performance. In particular, our work identifies the “low-temperature (local) MCMC threshold” for both sparse models, significantly improving upon a series of related prior works on the topic.
Notably, in the sparse tensor PCA model our results indicate that low-temperature local MCMC methods greatly underperform compared to other studied time-efficient methods, such as the class of low-degree polynomials. Moreover, the “amount” by which they underperform appears connected to the size of the computational-statistical gap of the model. Specifically, an intriguing mathematical pattern is observed in sparse tensor PCA, which we call “the geometric rule”: the information-theory threshold, the threshold for low-degree polynomials and the MCMC threshold form a geometric progression for all values of the sparsity level and tensor power.
@@ 1. Introduction
Over the recent years, it has been revealed that a plethora of statistical estimation models exhibit computational-statistical gaps, i.e., parameter regimes for which some estimator can achieve small error, but no such polynomial-time estimator is known to achieve similar error guarantees. It is an intriguing and rapidly growing research direction to understand whether these gaps are “inherent” or an inability of our current polynomial-time statistical estimation techniques. Given that the (simpler [^1] Often the statistical models of interest include random noise assumptions, hence cannot be treated as “classic” computational questions about tasks with a worst-case input, such as e.g., the traveling salesman problem over a worst-case graph. It is this difference that makes statements about the classes such as $\mathcal{P}$ and $\mathcal{NP}$ not directly applicable to these statistical estimation settings.) worst-case question of $\mathcal{P}\neq\mathcal{NP}$ remains open in computational complexity theory, unfortunately proving that all polynomial-time estimators fail to close any given computational-statistical gap appears well beyond our current mathematical abilities.
@@
Despite that, over the recent years, researchers have constructed several tools to provide “rigorous evidence” for a computational-statistical gap of interest. Some of the profound approaches include average-case reductions to “hard regimes” of other statistical problems (see e.g., [8, 46, 10]), statistical physics inspired “hardness” heuristics [25], and also proofs that natural restricted classes of polynomial-time estimators fail to close these gaps. Some of the most popular studied such restricted classes are MCMC methods [36, 16], low-degree polynomials [59], semidefinite programs [5], message passing procedures [23], and the class of statistical query methods [22].
@@
Among these restricted classes of estimators perhaps the most recent attention has been received by the class of estimators that take the form of a low-degree (e.g., $O(1)$ or $O(\log n)$ degree) polynomial function of the data (first used in the context of hypothesis testing in [35] and for estimation settings in [59]). A series of papers have built interesting tools to understand the exact threshold (in terms of some critical natural quantity of the model, often the sample size or signal to noise ratio) for a given statistical problem above which some low-degree polynomial works, and below which all low-degree polynomials fail (see e.g., [38] for a survey). On top of that, a conjecture has been put forward in the context of hypothesis testing [35] that – under mild noisy and symmetry assumptions [34, 69] – the threshold below which all $O(\log n)$-degree polynomials fail should in fact coincide with the threshold for which all polynomial-time methods fail as well. In other words, subject to this “low-degree conjecture”, in many classical statistical models the performance of $O(\log n)$-degree polynomials (which one can often directly analyze) allow us to characterize the performance of all polynomial-time estimators. Given the recent rapid growth of the low-degree literature for estimation problems following [59] (e.g., [50, 48, 43]), it is expected that a similar low-degree conjecture will be put forward in the near future in terms of statistical estimation problems as well.
@@
Despite this recent popularity of analyzing the performance of low-degree polynomials for statistical estimation settings, it is natural to wonder what happens with other natural classes of polynomial-time estimators that are significantly different from low-degree polynomials (e.g., estimators that are not even well-approximated by them). One such highly canonical class of statistical estimators (both in theory and applications of statistics) is the class of Markov Chain Monte Carlo (MCMC) methods that try to sample from the posterior (or slight variants of it), such as the Metropolis chain in discrete time and Langevin dynamics in continuous time. Researchers have studied this important class for a large family of models, but often arrive at significantly less satisfying theoretical results than the ones for the literature of low-degree methods. Among these MCMC results, most of them concern lower bounds, that is provable failures of classes of MCMC method to achieve a certain guarantee. Some influential such results include MCMC lower bounds for the planted clique problem [36, 28, 16], sparse principal component analysis (PCA) [24, 2] and tensor PCA [7]. Importantly, many of the recent such results leverage a suggested proof framework known as Overlap Gap Property for inference to obtain bottlenecks for the studied Markov chains, first introduced in [26]. On the other hand, there are significantly fewer positive MCMC results known for such estimation problems, including though interesting positive results about Bayesian variable selection [67], tensor PCA [7], and recently also the planted clique model [29] and the stochastic block model [42]. Yet, even concerning these relatively few positive results, most of them are an outcome of a technical detailed analysis tailored to the specific model of interest. In particular, as opposed to lower bounds and the proof framework of [26], there is a lack of a general “easy-to-use” strategy of how can one obtain a positive MCMC result for a new model of interest, besides directly trying to bound the mixing time (or analyze its locally stationary measures as in the recent work [42]). A significant motivation for this work is to answer the following question:
@@
Besides only obtaining positive MCMC results, it is of course highly desirable to obtain the “optimal” MCMC results, in the sense of locating exactly what is the minimal sample size (or signal to noise ratio) so that an element of some natural class of MCMC methods succeeds. Unfortunately, it is significantly understudied how one could locate exactly the “MCMC threshold” for a generic class of estimation settings, in sharp contrast with the generally applicable tools designed for the class of low-degree polynomials [59, 68]. On top of that, the pursuit of the MCMC threshold in estimation settings is even more motivated because of recent surprising discoveries in the context of tensor PCA [7] and planted clique [16]. The reason is that in both of these models the cited works have proven that a severe underperformance of certain natural MCMC methods takes place compared to other polynomial-time methods (something also noticed in the statistical physics literature [47]). To make it more precise, let us focus on the tensor PCA case where for some integer $t\geq 2$ one observes the $t$-tensor $Y=\lambda v^{\otimes t}+W$ for some $v\in S^{d-1}$ chosen uniformly at random from the sphere and $W$ has i.i.d. $N(0,1)$ entries. The task is to recover $v$ by observing $Y$. While the minimum $\lambda=\lambda_{d}>0$ for which this is possible by some (time-unconstrained) method scales like $\lambda_{\mathrm{STATS}}\approx\sqrt{d}$, polynomial-time methods are known to begin to work only when $\lambda_{\mathrm{ALG}}\approx d^{t/4}$ and below $\lambda_{\mathrm{ALG}}$ all $O(\log n)$-degree polynomials are known to fail, suggesting that $\lambda_{\mathrm{ALG}}$ is the computational threshold for this setting. Yet, if one runs Langevin dynamics that tries to sample from the (slightly temperature misparametrized) posterior, [7] proves that the dynamics can recover $v$ in polynomial-time at the threshold $\lambda_{\mathrm{MCMC}}\approx d^{t/2-1/2}$. Notice that for $t\geq 3$, $d^{t/4}=o(d^{t/2-1/2})$. It is perhaps striking the underperformance by a polynomial-in-$d$ factor of the Langevin dynamics for this seemingly simple task. For such reasons, this phenomenon that tensor PCA undergoes is known in the community as a “local-to-computational gap”. A partial motivation of this work is to understand how generic the existence of local-to-computational gaps is in statistical estimation settings.
@@
Besides the mathematical elegance, a positive answer to such a question would mean that for a class of models, as long as there is a computational statistical trade-off, i.e., $\lambda_{\mathrm{ALG}}/\lambda_{\mathrm{STATS}}=\omega(1)$, there is a local-to-computational gap as well, i.e., $\lambda_{\mathrm{MCMC}}/\lambda_{\mathrm{ALG}}=\omega(1)$. In this work, we interestingly reveal that the geometric rule does indeed hold beyond just the tensor PCA model, and in particular for the sparse matrix and tensor PCA models as well.
@@ 1.1. Contributions
We study the discrete parameter estimation set-up, where for some parameter $\theta^{*}\in\Theta\subseteq\mathbb{R}^{d}$ sampled from some prior $\mathcal{P}$, we observe a dataset $\mathcal{D}\sim\mathbb{P}_{\theta^{*}}$. We assume that $\Theta$ is a finite set of parameters and the prior is fully supported on $\Theta$. The goal of the statistician is to exactly recover the value of $\theta^{*}$ from observing $\mathcal{D}$, with high probability over both the randomness of the prior and the channel as $d$ grows to infinity. Here and everywhere in the paper, by saying that an event holds with high probability as $d$ grows to infinity, we mean that it’s probability converges to one as $d$ grows to infinity.
@@
For convenience in what follows, we refer to such a family of local Markov chains as “low-temperature MCMC methods”. We are interested in under which assumptions low-temperature MCMC methods can recover $\theta^{*}$ with high probability.
@@ 1.1.1. A general positive result
As we mentioned in the Introduction, the positive results for such low-temperature MCMC methods are few, often quite technical, and usually an outcome of an analysis crucially tailored to each case of interest.
Our first contribution is a general positive “black-box” result (Theorem 2.3) which offers a structural geometric condition under which some low-temperature MCMC method succeeds in polynomial-time to optimize the posterior, and in particular recover $\theta^{*}$ with high probability. Our result in fact applies for any optimization problem over a discrete feasible set and we expect to be of independent interest to the discrete optimization community as well.
@@
As long as solely this condition is satisfied, we prove that simply running the (low-temperature) Metropolis chain on this neighborhood graph for polynomial time is guaranteed to find the optimizer of the posterior $\theta^{*}$ with high probability. As we mentioned above, to the best of our knowledge, no similar result is known at that level of generality even in the discrete/combinatorial optimization community. The proof is simple and is based on a general canonical paths construction [37] and could be potentially of independent interest across communities.
To show the generality of Theorem 2.3 we derive two positive MCMC results for two of the most classical, but significantly different, sparse estimation problems with well-studied computational-statistical gaps. First, we focus on the sparse tensor PCA model (also called, constant high order clustering) [44, 17] which is a special case of a Gaussian additive model, i.e., one observes a sparse signal drawn from some prior where Gaussian noise is added to it. Second, we also apply it to Gaussian sparse regression [26] which is a special case of a Linear Model, i.e., one observes i.i.d. samples of a (random) linear transformation of a sparse signal where Gaussian noise is added to each sample. Interestingly, using the Overlap Gap Property framework from [26] we can also show that both the obtained low-temperature MCMC results are tight, in the sense that a large class of low-temperature MCMC results fail to work in polynomial-time, whenever the conditions of our positive result do not hold. In particular, these results allow us to characterize the exact power of a large class of low-temperature MCMC methods for both of these settings. We finally highlight that both applications of the Overlap Gap Property framework are rather non-trivial and require highly technical applications of the second moment method to prove existence of “bad” bottlenecks in the landscape of the models.
@@ 1.1.2. Sparse Tensor PCA
The sparse tensor PCA model we focus on is a canonical Gaussian principal component analysis setting with a large literature in theoretical statistics (see e.g., [12, 13, 14, 66, 53, 44] and references therein). Moreover, the setting is greatly motivated by a wide range of applications in many domains, such as genetics, social sciences, and engineering (see e.g., [33] for a survey). More specific interesting examples include topics such as microbiome studies [21] and multi-tissue multi-individual gene expression data [65].
Now, from the theoretical standpoint it is known [44, 53] that the optimal estimator (MAP) can recover $v$ if and only if $\lambda\geq\lambda_{\mathrm{STATS}}(t,k)$ for some
@@
Moreover, based on average-case reductions to hypergraphs variants of the planted clique problem [10, 44], polynomial-time estimators are expected to work if and only if $\lambda\geq\lambda_{\mathrm{ALG}}(t,k)$ for some
@@
Note that for all $k=o(p),t\geq 2$ it holds $\lambda_{\mathrm{ALG}}(t,k)=\tilde{\omega}(\lambda_{\mathrm{STATS}}(t,k))$ and therefore the model is conjectured to exhibit a computational-statistical trade-off.
@@
It should be noted that the mentioned algorithmic lower bound when $\lambda=o(\lambda_{\mathrm{ALG}}(t,k))$ from [10, 44] applies in principle for a slight variant of the model of sparse tensor PCA compared to one we introduced above, which the authors of [44] denote by $\mathrm{CHC}_{R}$. In $\mathrm{CHC}_{R}$ one observes $Y=\lambda v_{1}\otimes v_{2}\otimes\ldots\otimes v_{t}+W$ for $v_{i},i=1,\ldots,t$ which are independent $k$-sparse binary vectors chosen uniformly at random. Note that in our case we focus instead on the case $v_{1}=\ldots=v_{t}=v$ where $v$ is a $k$-sparse binary vectors chosen uniformly at random. Often in the literature, the model we study is called the symmetric version of a tensor PCA model, while the one studied in [44] is called the asymmetric version (see e.g., the discussion in [19]).
Yet, a quick check in the proof of their lower bound [44, Theorem 16] shows that in fact the authors first prove the conditional failure of polynomial-time algorithms when $\lambda=o(\lambda_{\mathrm{ALG}}(t,k))$ for our case of interest, that is when $v_{1}=\ldots=v_{t}$, and then add an extra step to also prove it for the case that $v_{i},i=1,\ldots,t$ are independent. In particular, their result applies almost identically to the case of symmetric sparse tensor PCA we focus on this work. For reasons of completeness, in Section 11, we describe exactly the slightly modified argument from [44] and how it applies in our symmetric setting. We finally also note that the reduction argument for sparse tensor PCA in [44] is based on some instrumental techniques introduced in [10] for the hardness of tensor PCA in the special case of $v\in\{-1,1\}^{n}$ (that is the case $v$ is Rademacher-valued and not sparse).
A careful application of our “black box” positive MCMC result interestingly implies that for all $t\geq 2$ and various values of the sparsity level $\omega(1)=k=o(p),$ if $\beta>0$ is large enough, the Metropolis chain on the Johnson graph with stationary measure $\pi_{\beta}$ hits the planted vector $v$ in polynomial time as long as $\lambda\geq\lambda_{\mathrm{MCMC}}(t,k)$ for a significantly different threshold compared to $\lambda_{\mathrm{ALG}}(t,k)$ given by
@@
See Theorem 3.3 for the exact statement and specifically the Johnson adjacency graph the Metropolis chain is naturally defined on. Note also that for all $k=o(p),t\geq 2$ it holds $\lambda_{\mathrm{MCMC}}(t,k)=\tilde{\omega}(\lambda_{\mathrm{ALG}}(t,k))$.
Satisfyingly, we also prove that when $\lambda=\tilde{o}(\lambda_{\mathrm{MCMC}}(t,k))$ a negative result for all low-temperature reversible local MCMC methods (including the Metropolis chain) holds. Specifically, we prove in Theorem 3.8 that for large enough $\beta>0$ all local reversible Markov chains with stationary measure $\pi_{\beta}$ require super-polynomial time to hit $v$. Notably, our result significantly improves upon the prior work [2] on sparse PCA where a similar negative result for local Markov chains was proven but only when $t=2$ and importantly under the more restrictive condition $\lambda=o(\lambda_{\mathrm{ALG}}(2,k))$. The proof of Theorem 3.8 is the most technical contribution of the present work. We prove – via a delicate conditional second moment method – a bottleneck for the MCMC dynamics based on the Overlap Gap Property framework from [26]. The proof carefully leverages the “flatness” idea from random graph theory [3, 28], but adjusted to high dimensional Gaussian tensors, and could be of independent interest. Moreover, the fact that the Overlap Gap Property from [26] appears in the landscape of the problem well into it’s “easy” phase where some polynomial-time method works (specifically when $\tilde{\omega}(\lambda_{\mathrm{ALG}}(t,k))=\lambda=\tilde{o}(\lambda_{\mathrm{MCMC}}(t,k))$) is a surprising fact, first conjectured in [28] for the planted clique model that remains unproved. This is a result which in fact aligns with similar results in the general realm of average-case complexity (see e.g., [41] for a similar recent result for average-case optimization) and naturally leads to multiple future questions about the relation between “geometric” complexity and algorithmic hardness.
@@
We remark that this is a significant departure compared to the established tensor PCA result [7] as described in (1.1), as it holds firstly under the sparsity assumption, and also it importantly applies even to the matrix case $t=2$ where tensor PCA does not exhibit either a computational-statistical gap or local-to-computational gap. Albeit interesting in its own right, the fact that the result applies when $t=2$ allows us also to see the geometric rule in simulations for relatively small values of $p,k$, which we find particularly satisfying (see Section 6). We suspect that a generalization of the geometric rule to a large family of Gaussian additive models is possible. The exact reason such an elegant formula appears in that generality remains an intriguing question for future work.